HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

A NUMERICAL-TENSORIAL “HYBRID” NUCLEAR-MOTION HAMILTONIAN AND DIPOLE MOMENT FOR SPECTRA CALCULATION

OF POLYATOMIC NONRIGID MOLECULES

\underline{MICHAËL REY} , \underline{DOMINIKA VIGLASKA} , Groupe de Spectrométrie Moléculaire et Atmosphérique UMR CNRS 7331, UFR Sciences BP 1039, 51687 Reims Cedex 2, France; OLEG EGOROV, ANDREI V. NIKITIN, Laboratory of

Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Faculty of Physics Tomsk State

University 36, Lenin Ave., Tomsk, 634050 Russia

The analysis and modelling of high-resolution spectra of nonrigid molecules require a specific Hamiltonian and group-theoretical formulation which differs significantly from that of more familiar rigid systems. Within the framework of the Hougen-Bunker-Johns (HBJ) theory^{1}, this work is devoted to the construction of a nonrigid Hamiltonian based on a suitable combination of numerical calculations for the nonrigid part in conjunction with the irreducible tensor operator method for the rigid part. A so-called “hybrid” model is thus introduced^{2}. For the first time, a variational calculation from ab initio potential energy surfaces is performed using the hybrid formulation of the exact HBJ kinetic energy operator expressed in terms of vibrational, large-amplitude motion and rotational tensor operators through the use of both curvilinear and rectilinear coordinates. Group theory for nonrigid molecules plays a central role in the characterization of the overall tunnelling splittings and will be discussed. The construction of the dipole moment operator will be also examined.

Validation tests consisting in a careful convergence study of the energy levels as well as in a comparison of results obtained from independent computer codes are given for the nonrigid molecules CH_{2}, CH_{3}, NH_{3} and H_{2}O_{2}. This work−combined with our recent method for constructing effective models^{3}−will pave the way to the analysis of high-resolution spectra of larger nonrigid systems.

This work is supported by the french Agence Nationale de la Recherche (ANR TEMMEX, Grant No. 21-CE30-0053-01) and the Russian Scientific Foundation (RSF, No. 22-42-09022).

1J. T. Hougen, P. R. Bunker, J. W. C. Johns, J. Mol. Spectrosc., 34 (1970) 136

2M. Rey, D. Viglaska, O. Egorov, A. V. Nikitin, (2023, submitted).

3M. Rey, J. Chem. Phys. 156 (2022) 224103. doi.org/10.1063/5.0089097 p-number: p091  Submitted on Mon, 12 Jun 2023 19:59:56 +0200